# and knot coloring invariants

### v2

@@ 3.1. GL-racks
Virtual racks are algebraic structures that can distinguish framed links in certain lens spaces and framed virtual links in thickened surfaces; see, for example, [virtual]*Section 3.2.
By Definition 3.1, GL-racks are precisely virtual racks in which all inner automorphisms $s_{x}$ are endomorphisms of virtual racks. Equivalently, a GL-rack $(X,s,\mathtt{u})$ is a virtual rack in which the operator group of $(X,s)$ identifies $x$ with $\mathtt{u}(x)$ for all $x\in X$; see [fenn]*Section 1.1.
@@
Example 3.5.
[bi]*Example 3.7 Given a permutation rack $P=(X,\sigma)_{\mathrm{perm}}$, a GL-structure on $P$ is precisely a permutation $\mathtt{u}\in S_{X}$ such that $\mathtt{u}\sigma=\sigma\mathtt{u}$. Given such a $\mathtt{u}$, we say that $(P,\mathtt{u})$ is a permutation GL-rack or constant action GL-rack, and we denote it by $(X,\sigma,\mathtt{u})_{\mathrm{perm}}$.
Example 3.6.
[bi]*Example 3.6 Let $G$ be a group, let $z\in Z(G)$ be a central element of $G$, and define $f:G\rightarrow G$ by $g\mapsto zg$. Then $(\operatorname{Conj}G,f)$ is a GL-quandle.
@@ 3.1.1. Bi-Legendrian racks
Next, we reproduce the definition of GL-racks given in the literature; see [karmakar]*Definition 3.1 or [bi]*Definition 3.2. Following Kimura [bi], we temporarily use the term bi-Legendrian racks to distinguish them from GL-racks in the sense of Definition 3.1. We also define Legendrian racks, which Ceniceros et al. [ceniceros] introduced in 2021.
Definition 3.9.
[karmakar]*Definition 3.1 Given a rack $R=(X,s)$, a bi-Legendrian structure on $R$ is a pair $(\mathtt{u},\mathtt{d})$ of maps $\mathtt{u},\mathtt{d}:X\to X$ satisfying the following axioms for all elements $x\in X$:
@@ 3.2. Equivalence of definitions
In light of these results, we will henceforth call bi-Legendrian racks GL-racks except when specifically citing the axioms in Definition 3.9 or denoting GL-racks as quadruples with $\mathtt{d}$. The previous two propositions make computing and classifying GL-racks significantly easier; see Section 4 and Appendix A.2. As a bonus, they also yield the following converse of [ceniceros]*Remark 2.
@@ 3.3.2. Free GL-racks
As with racks, the free GL-rack $\operatorname{Free}_{\mathsf{GLR}}(X)$ on a set $X$ exists and is uniquely (up to isomorphism) characterized by a universal property; see [karmakar]*Proposition 4.2 for a detailed derivation. Namely, there exists a set map $i:X\to\operatorname{Free}_{\mathsf{GLR}}(X)$ such that for all GL-racks $R=(Y,t,\mathtt{u})$ and all set maps $\varphi_{0}:X\to Y$, there exists a unique GL-rack homomorphism $\varphi\in\operatorname{Hom}_{\mathsf{GLR}}(\operatorname{Free}_{\mathsf{GLR}}(X),R)$ such that $\varphi i=\varphi_{0}$. That is, the following diagram commutes:
@@
Karmakar et al. [karmakar]*Section 4 gave an explicit set-theoretic construction of $\operatorname{Free}_{\mathsf{GLR}}(X)$. Here, we simplify their construction using Definition 3.1 and and Proposition 3.12.
Definition 3.15.
[karmakar]*Section 4 Let $X$ be a set. We define the free GL-rack on $X$ as follows. If $X=\emptyset$, let $\operatorname{Free}_{\mathsf{GLR}}(X)$ be the empty GL-rack. Else, let the universe of words generated by $X$ be the set $W(X)$ such that $X\subset W(X)$ and $s_{y}(x),\,s_{y}^{-1}(x),\,\mathtt{u}(x),\,\mathtt{u}^{-1}(x)\in W(X)$ for all $x,y\in W(X)$. Let $V(X)$ be the set of equivalence classes of elements of $W(X)$ modulo the congruence generated by the following relations for all $x,y,z\in W(X)$:
@@ 3.3.3. Free objects in subtheories of GL-racks
Since $\mathsf{GLR}$ is the category of models of the algebraic theory of GL-racks in $\mathsf{Set}$, the free GL-rack $L=\operatorname{Free}_{\mathsf{GLR}}(X)$ on the one-element set $X=\{x\}$ is a strong generator or separator for $\mathsf{GLR}$; see, for example, [alg-theories]*Proposition 3.3.3.
@@ 3.3.5. Mediality and Hom GL-racks
4. Classification of GL-structures, automorphisms, and centers
In 2024, Karmakar et al. [karmakar]*Section 3 posed the following question: what are all the possible GL-structures on a given rack? In this section, we answer this question and classify various infinite families of GL-racks. As applications, we discuss automorphism groups of GL-racks and compute these groups for all GL-racks whose underlying racks are dihedral quandles. We also compute the centers of $\mathsf{GLR}$, $\mathsf{GLQ}$, and the categories of Legendrian racks and Legendrian quandles.
@@ 4.2.2. Conjugation GL-quandles
If $G$ is abelian, then $Q$ is a trivial quandle, so the first claim follows from Example 4.4. Next, recall a result of Elhamdadi et al. [dihedral]*Theorem 2.3 that $\operatorname{Inn}Q=\operatorname{Inn}_{\mathsf{Grp}}G$ for all groups $G$. Also, recall a result of Bardakov et al. [centerless]*Corollary 2 that $\operatorname{Aut}Q=\operatorname{Aut}_{\mathsf{Grp}}G$ if and only if $G$ is centerless. In this case, Theorem 4.1 states that
@@ 4.2.3. Takasaki GL-kei
We first recall a classification result of Bardakov et al. [takasaki-aut]*Theorem 4.2. For all abelian additive groups $A$ without $2$-torsion, $\operatorname{Aut}T(A)$ is isomorphic to the holomorph
@@ 4.2.4. Dihedral GL-quandles
Without the assumption that $A$ is $2$-torsion-free, there are infinitely many counterexamples to Proposition 4.9. To show this, we will complete the classification of GL-structures on dihedral quandles $R_{n}$ using the following results of Elhamdadi et al. [dihedral]*Theorems 2.1 and 2.2. The automorphism group $\operatorname{Aut}R_{n}$ is the affine group of $\mathbb{Z}/n\mathbb{Z}$. Thus, $\operatorname{Aut}R_{n}$ is isomorphic to the holomorph
@@ 4.3.1. Automorphisms of dihedral GL-quandles
We classify automorphism groups of GL-racks whose underlying racks are dihedral quandles $R_{n}$ of order $n$. Once again, we use a result of Elhamdadi et al. [dihedral]*Theorem 2.1 to identify $\operatorname{Aut}R_{n}\cong\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$.
@@
4.4. The center of the category of GL-racks
Recall that the center of a category $\mathcal{C}$ is the commutative monoid $Z(\mathcal{C})$ of natural endomorphisms of the identity functor $\mathbf{1}_{\mathcal{C}}$. In 2018, Szymik [center]*Theorems 5.4 and 5.5 computed that $Z(\mathsf{Rack})=\langle\Theta\rangle\cong\mathbb{Z}$ and $Z(\mathsf{Qnd})\cong\{1\}$. In this subsection, we similarly compute the centers of $\mathsf{GLR}$ and various full subcategories of $\mathsf{GLR}$.
@@ 4.4. The center of the category of GL-racks
5. Categorical equivalence of racks and GL-quandles
In this section, we show that the categories of racks and GL-quandles are isomorphic in a way that preserves mediality. This surprising result generalizes the one-to-one correspondences observed in Appendix A.1 and induces isomorphisms of the respective algebraic theories.
5.1. Construction of $F$
@@ 5.3. Isomorphism of categories
Since $F$ and $G$ are left and right adjoints to each other, they both preserve limits and colimits. In particular, they preserve finite products and filtered colimits, so by [alg-theories]*Lemma 3.8.3, $F$ and $G$ are algebraic functors. Since $F$ and $G$ are equivalences of categories, the claim follows directly from [alg-theories]*Proposition 3.12.1. ∎
@@ 6. Tensor products of racks and GL-racks
In 2014, Crans and Nelson [Hom]*Section 8.1 categorified the results of Corollary 2.22 by considering universal-algebraic tensor products of medial quandles. However, these tensors remain unexplored in the literature.
@@ 6.1. Construction and universal property
We note that the medialized associated quandles of the tensor products of racks in Definition 6.1 recover the tensor products of medial quandles that Crans and Nelson [Hom]*Section 8.1 introduced in 2014. On the other hand, these tensors are distinct from the tensor products that Kamada [tensors]*Definition 3.1 introduced in 2021, which are sets without canonical rack structures.
@@ 6.1.1. Bihomomorphisms
The discussion in [alg-theories]*p. 171 shows that the tensor products in Definition 6.1 are each characterized by a universal factorizing property. Before we can state this property, we will need the following definition from universal algebra. This is a special case of a more general definition for arbitrary algebraic theories; see, for example, [alg-theories]*Definition 3.10.2.
@@ 6.1.2. Universal property of tensor products
This is a consequence of universal algebra. Regard racks and GL-racks as algebraic theories with binary operations $s_{-}(-)$ and $s^{-1}_{-}(-)$, along with unary operations $\mathtt{u}$ and $\mathtt{u}^{-1}$ for GL-racks. Due to Remark 6.2, the tensor product in Definition 6.1 is precisely the tensor product constructed in the proof of [alg-theories]*Theorem 3.10.3, and $\psi(x,y)=x\otimes y$ for all $(x,y)\in X\times Y$. Thus, the claim follows from the discussion in [alg-theories]*p. 171. ∎
@@ 6.1.3. Internal hom-tensor adjunctions
Mirroring the work of Crans and Nelson [Hom]*Theorem 12 on the category of medial quandles, the next result continues the analogy with the category of modules over a ring by describing internal hom-tensor adjunctions in $\mathsf{Rack}_{\mathrm{med}}$ and $\mathsf{GLR}_{\mathrm{med}}$.
@@
Recall that the algebraic theories of medial racks and medial GL-racks are commutative. Therefore, both claims are special cases of a general result for commutative algebraic theories; see, for example, [alg-theories]*Theorem 3.10.3.
In particular, $\otimes_{\mathrm{med}}$ is precisely the tensor product constructed in the proof of [alg-theories]*Theorem 3.10.3, and the tensor unit is the free rack (resp. GL-rack) $L$ on one element. Indeed, combining Example 2.19 (resp. Proposition 3.19) with Example 2.12 shows that $L$ is medial. ∎
@@ 6.2. Nonmedial tensor products
One may ask how much of the structure in Theorem 6.7 remains if we drop the mediality assumption. By [alg-theories]*Theorem 3.10.3, we lose the closed structure because racks and GL-racks are noncommutative algebraic theories. Nevertheless, we show that tensor products in $\mathsf{Rack}$ and $\mathsf{GLR}$ surprisingly retain tensor units.

### v3

@@ 3.1. GL-racks
Virtual racks are algebraic structures that can distinguish framed links in certain lens spaces and framed virtual links in thickened surfaces; see, for example, [virtual]*Sec. 3.2.
By Definition 3.1, GL-racks are precisely virtual racks in which all inner automorphisms $s_{x}$ are endomorphisms of virtual racks. Equivalently, a GL-rack $(X,s,\mathtt{u})$ is a virtual rack in which the operator group of $(X,s)$ identifies $x$ with $\mathtt{u}(x)$ for all $x\in X$; see [fenn]*Sec. 1.1.
@@
Example 3.5 ([bi]*Ex. 3.7).
Given a permutation rack $P=(X,\sigma)_{\mathrm{perm}}$, a GL-structure on $P$ is precisely a permutation $\mathtt{u}\in S_{X}$ such that $\mathtt{u}\sigma=\sigma\mathtt{u}$. Given such a $\mathtt{u}$, we say that $(P,\mathtt{u})$ is a permutation GL-rack or constant action GL-rack, and we denote it by $(X,\sigma,\mathtt{u})_{\mathrm{perm}}$.
Example 3.6 ([bi]*Ex. 3.6).
Let $G$ be a group, let $z\in Z(G)$ be a central element of $G$, and define $f:G\rightarrow G$ by $g\mapsto zg$. Then $(\operatorname{Conj}G,f)$ is a GL-quandle.
@@ 3.1.1. Bi-Legendrian racks
Next, we reproduce the definition of GL-racks given in the literature; see [karmakar]*Def. 3.1 or [bi]*Def. 3.2. Following Kimura [bi], we temporarily use the term bi-Legendrian racks to distinguish them from GL-racks in the sense of Definition 3.1. We also define Legendrian racks, which Ceniceros et al. [ceniceros] introduced in 2021.
Definition 3.9 ([karmakar]*Def. 3.1).
Given a rack $R=(X,s)$, a bi-Legendrian structure on $R$ is a pair $(\mathtt{u},\mathtt{d})$ of maps $\mathtt{u},\mathtt{d}:X\to X$ satisfying the following axioms for all elements $x\in X$:
@@ 3.2. Equivalence of definitions
In light of these results, we will henceforth call bi-Legendrian racks GL-racks except when specifically citing the axioms in Definition 3.9 or denoting GL-racks as quadruples with $\mathtt{d}$. The previous two propositions make computing and classifying GL-racks significantly easier; see Section 4 and Appendix A.2. As a bonus, they also yield the following converse of [ceniceros]*Rem. 2.
@@ 3.3.2. Free GL-racks
As with racks, the free GL-rack $\operatorname{Free}_{\mathsf{GLR}}(X)$ on a set $X$ exists and is uniquely (up to isomorphism) characterized by a universal property; see [karmakar]*Prop. 4.2 for a detailed derivation. Namely, there exists a set map $i:X\to\operatorname{Free}_{\mathsf{GLR}}(X)$ such that for all GL-racks $R=(Y,t,\mathtt{u})$ and all set maps $\varphi_{0}:X\to Y$, there exists a unique GL-rack homomorphism $\varphi\in\operatorname{Hom}_{\mathsf{GLR}}(\operatorname{Free}_{\mathsf{GLR}}(X),R)$ such that $\varphi i=\varphi_{0}$. That is, the following diagram commutes:
@@
Karmakar et al. [karmakar]*Sec. 4 gave an explicit set-theoretic construction of $\operatorname{Free}_{\mathsf{GLR}}(X)$. Here, we simplify their construction using Definition 3.1 and and Proposition 3.12.
Definition 3.15 ([karmakar]*Sec. 4).
Let $X$ be a set. We define the free GL-rack on $X$ as follows. If $X=\emptyset$, let $\operatorname{Free}_{\mathsf{GLR}}(X)$ be the empty GL-rack. Else, let the universe of words generated by $X$ be the set $W(X)$ such that $X\subset W(X)$ and $s_{y}(x),\,s_{y}^{-1}(x),\,\mathtt{u}(x),\,\mathtt{u}^{-1}(x)\in W(X)$ for all $x,y\in W(X)$. Let $V(X)$ be the set of equivalence classes of elements of $W(X)$ modulo the congruence generated by the following relations for all $x,y,z\in W(X)$:
@@ 3.3.3. Free objects in subtheories of GL-racks
Since $\mathsf{GLR}$ is the category of models of the algebraic theory of GL-racks in $\mathsf{Set}$, the free GL-rack $L=\operatorname{Free}_{\mathsf{GLR}}(X)$ on the one-element set $X=\{x\}$ is a strong generator or separator for $\mathsf{GLR}$; see, for example, [alg-theories]*Prop. 3.3.3.
@@ 3.3.5. Mediality and Hom GL-racks
[…]
Corollary 3.21.
Let $\mathcal{G}(\Lambda)$ be the fundamental GL-rack of a Legendrian link $\Lambda$ (see [karmakar]). Then for all medial GL-racks $M$, the isomorphism type of $\operatorname{Hom}_{\mathsf{GLR}}(\mathcal{G}(\Lambda),M)$ as a medial GL-rack is an invariant of $\Lambda$.
4. Group-theoretic aspects of GL-racks
In 2024, Karmakar et al. [karmakar]*Sec. 3 posed the following question: what are all the possible GL-structures on a given rack? In this section, we answer this question and classify various infinite families of GL-racks. As applications, we discuss automorphism groups of GL-racks and compute these groups for all GL-racks whose underlying racks are dihedral quandles.
@@ 4.2.2. Conjugation GL-quandles
If $G$ is abelian, then $Q$ is a trivial quandle, so the first claim follows from Example 4.4. Next, recall a result of Elhamdadi et al. [dihedral]*Thm. 2.3 that $\operatorname{Inn}Q=\operatorname{Inn}_{\mathsf{Grp}}G$ for all groups $G$. Also, recall a result of Bardakov et al. [centerless]*Cor. 2 that $\operatorname{Aut}Q=\operatorname{Aut}_{\mathsf{Grp}}G$ if and only if $G$ is centerless. In this case, Theorem 4.1 states that
@@ 4.2.3. Takasaki GL-kei
We first recall a classification result of Bardakov et al. [takasaki-aut]*Thm. 4.2. For all abelian additive groups $A$ without $2$-torsion, $\operatorname{Aut}T(A)$ is isomorphic to the holomorph
@@ 4.2.4. Dihedral GL-quandles
Without the assumption that $A$ is $2$-torsion-free, there are infinitely many counterexamples to Proposition 4.9. To show this, we will complete the classification of GL-structures on dihedral quandles $R_{n}$ using the following results of Elhamdadi et al. [dihedral]*Thms. 2.1 and 2.2. The automorphism group $\operatorname{Aut}R_{n}$ is the affine group of $\mathbb{Z}/n\mathbb{Z}$. Thus, $\operatorname{Aut}R_{n}$ is isomorphic to the holomorph
@@ 4.3.1. Automorphisms of dihedral GL-quandles
We classify automorphism groups of GL-racks whose underlying racks are dihedral quandles $R_{n}$ of order $n$. Once again, we use a result of Elhamdadi et al. [dihedral]*Thm. 2.1 to identify $\operatorname{Aut}R_{n}\cong\mathbb{Z}/n\mathbb{Z}\rtimes(\mathbb{Z}/n\mathbb{Z})^{\times}$.
@@
5. Categorical aspects of GL-racks
In this section, we compute the centers of $\mathsf{GLR}$ and several of its full subcategories. We also prove a surprising equivalence of categories between $\mathsf{Rack}$ and $\mathsf{GLQ}$.
5.1. Categorical centers
Recall that the center of a category $\mathcal{C}$ is the commutative monoid $Z(\mathcal{C})$ of natural endomorphisms of the identity functor $\mathbf{1}_{\mathcal{C}}$. In 2018, Szymik [center]*Thms. 5.4 and 5.5 computed that $Z(\mathsf{Rack})=\langle\Theta\rangle\cong\mathbb{Z}$ and $Z(\mathsf{Qnd})\cong\{1\}$. In this section, we similarly compute the centers of $\mathsf{GLR}$, $\mathsf{GLQ}$, and the categories of Legendrian racks and Legendrian quandles.
@@ 5.1. Categorical centers
5.2. Categorical equivalence of racks and GL-quandles
Next, we show that the categories of racks and GL-quandles are isomorphic in a way that preserves mediality. This surprising result generalizes the one-to-one correspondences observed in Appendix A.1 and corresponds to isomorphisms of the respective algebraic theories.
5.2.1. Construction of $F$
@@ 5.2.3. Isomorphism of categories
Since $F$ and $G$ are left and right adjoints to each other, they both preserve limits and colimits. In particular, they preserve finite products and filtered colimits, so by [alg-theories]*Lem. 3.8.3, $F$ and $G$ are algebraic functors. Since $F$ and $G$ are equivalences of categories, the claim follows directly from [alg-theories]*Prop. 3.12.1. ∎
@@ 6. Tensor products of racks and GL-racks
In 2014, Crans and Nelson [Hom]*Sec. 8.1 categorified the results of Corollary 2.22 by considering universal-algebraic tensor products of medial quandles. However, these tensors remain unexplored in the literature.
@@ 6.1. Construction and universal property
We note that the medialized associated quandles of the tensor products of racks in Definition 6.1 recover the tensor products of medial quandles that Crans and Nelson [Hom]*Sec. 8.1 introduced in 2014. On the other hand, these tensors are distinct from the tensor products that Kamada [tensors]*Def. 3.1 introduced in 2021, which are sets without canonical rack structures.
@@ 6.1.1. Bihomomorphisms
The discussion in [alg-theories]*p. 171 shows that the tensor products in Definition 6.1 are each characterized by a universal factorizing property. Before we can state this property, we will need the following definition from universal algebra. This is a special case of a more general definition for arbitrary algebraic theories; see, for example, [alg-theories]*Def. 3.10.2.
@@ 6.1.2. Universal property of tensor products
This is a consequence of universal algebra. Regard racks and GL-racks as algebraic theories with binary operations $s_{-}(-)$ and $s^{-1}_{-}(-)$, along with unary operations $\mathtt{u}$ and $\mathtt{u}^{-1}$ for GL-racks. Due to Remark 6.2, the tensor product in Definition 6.1 is precisely the tensor product constructed in the proof of [alg-theories]*Thm. 3.10.3, and $\psi(x,y)=x\otimes y$ for all $(x,y)\in X\times Y$. Thus, the claim follows from the discussion in [alg-theories]*p. 171. ∎
@@ 6.1.3. Internal hom-tensor adjunctions
Mirroring the work of Crans and Nelson [Hom]*Thm. 12 on the category of medial quandles, the next result continues the analogy with the category of modules over a ring by describing internal hom-tensor adjunctions in $\mathsf{Rack}_{\mathrm{med}}$ and $\mathsf{GLR}_{\mathrm{med}}$.
@@
Recall that the algebraic theories of medial racks and medial GL-racks are commutative. Therefore, both claims are special cases of a general result for commutative algebraic theories; see, for example, [alg-theories]*Thm. 3.10.3.
In particular, $\otimes_{\mathrm{med}}$ is precisely the tensor product constructed in the proof of [alg-theories]*Thm. 3.10.3, and the tensor unit is the free rack (resp. GL-rack) $L$ on one element. Indeed, combining Example 2.19 (resp. Proposition 3.19) with Example 2.12 shows that $L$ is medial. ∎
@@ 6.2. Nonmedial tensor products
One may ask how much of the structure in Theorem 6.7 remains if we drop the mediality assumption. By [alg-theories]*Thm. 3.10.3, we lose the closed structure because racks and GL-racks are noncommutative algebraic theories. Nevertheless, we show that tensor products in $\mathsf{Rack}$ and $\mathsf{GLR}$ surprisingly retain tensor units.